OpenStudy (anonymous):

need help solving \[(\sqrt{11}+2)(2\sqrt{11}-1)\]

OpenStudy (anonymous):

solve this multiplication \[\sqrt{11}\times 2\sqrt{11}+2\times2\sqrt{11}-\sqrt{11}-2\] \[=2\times11+4\sqrt{11}-\sqrt{11}-2\]

OpenStudy (anonymous):

\[3\sqrt{11}+20\]

OpenStudy (anonymous):

thanks